I need to know the real and imaginary parts of the complex number.
(5+i)/(2-i)

Answer by venugopalramana(3286)

I need to know the real and imaginary parts of the complex number.
(5+i)/(2-i)...MULTIPLY AND DIVIDE WITH CONJUGATE OF D.R.
=(5+i)(2+i)/(2-i)(2+i)
=(10+5i+2i+i^2)/(4-i^2)
=(9+7i)/5=(9/5)+(7/5)i
R.P.=9/5
I.P.=7/5